OpenStudy (anonymous):

Line CD contains points A (4, 6) and B (−2, 6). The slope of line CD is zero undefined positive negative

OpenStudy (anonymous):

\[m = \frac{ y_2-y_1 }{ x_2-x_1 }\] is your slope if you're given a line with the points (4,6) and (-2,6) you can just plug it in the formula provided above \[m = \frac{ 6-6 }{ -2-4 } = ...?\] any idea what's going to happen?

OpenStudy (anonymous):

would it be undefined?

OpenStudy (usukidoll):

if we have something that's undefined it means that the denominator is 0 and we don't want that. We can't divide by 0 1/0 <-- this doesn't make sense. 0/1 < --- makes sense because 0 divided by 1 is 0

OpenStudy (anonymous):

 not quite, you would get a horizontal line, when you have 0 in the numerator your answer is 0, when you have 0 in the denominator you would then get undefined, so what's the case here?

OpenStudy (anonymous):

A?

OpenStudy (usukidoll):

\[m = \frac{ 6-6 }{ -2-4 } = ...? \] what's 6-6 what's -2-4